Perform the indicated operations

17/90 + 13/60 -11/50
The answers I came up with have all been wrong. 7/25,3/50/563/900,0.185

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

17/90 + 13/60 -11/50
-----------
Find a common denominator.
900 works
---
17/90 + 13/60 -11/50 = (170 + 195 - 198)/900
= 167/900
= 0.18555...

Answer by stanbon(75887)

Perform the indicated operations
17/90 + 13/60 -11/50
------
lcd = 900
---
= [170 + 195 - 198]/900
----
= 167/900
-----
= 0.1855...
Cheers,
Stan H.
=================